OpenStudy (anonymous):

counterexample of 9^n -1 is divisble by 8

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 9^{~n-1} }\) by 8 ?

OpenStudy (solomonzelman):

when I see the power of n-1 I think of derivatives.

OpenStudy (solomonzelman):

What do you think ?

OpenStudy (zzr0ck3r):

I am guessing its \(9^n-1\)

OpenStudy (solomonzelman):

I would say either way it is ....

OpenStudy (solomonzelman):

Not telling the answer straight out

OpenStudy (zzr0ck3r):

yeah....

OpenStudy (anonymous):

The first two steps will be provided. For step 3, match the right and left sides of the equation to prove the statement. Step 1: When n = 1, 9n -1 = 91 - 1 = 8. Eight is divisible by 8, so the equation is true for n = 1 Step 2: Assume that 9k-1 is divisible by 8 for some positive integer, k. This means that there is a whole number r such that 9k-1 = 8r Step 3: Show that the statement n = k + 1 is true 9^k -1 9^k 9(9^k) 9^k+1 9^k-1 -1 answers: 8r 8r+1 72r+9 9(8r+1) 8(9r+1) 72r+8